Let f(x) = 3x2 + 4 and  g(x) = 2x − 4.

Find the function.

(gf)(x) =
Answer #1

ut, tex) a 3x7 y gex) = 2x-4 .. (gof)(x) g (Hex)) g(3x+4) = 2(3x+4) – 4 6278-4 = 6xt 4 (gof) (2) baty